OpenStudy (anonymous):

3*5^2*11 in standard form

OpenStudy (jack1):

hi @kate_0123 welcome to OS

OpenStudy (jack1):

3*5^2*11 = \[3 \times 5^2 \times 11 = 33 \times 5^2 = 33 \times 25 = ???\] @kate_0123 ???